# U-index: A Universal Indexing Framework for Matching Long Patterns

### v1

@@ 4 A Universal Indexing Framework for Matching Long Patterns
In this section, we describe a universal indexing framework for a text $T$ of length $n$ — referred to as the U-index — to retrieve all occurrences of a pattern $P$ of length at least $\ell$ in $T$. Refer to Figure 1 for an overview.
@@
Overview. The core idea of the U-index is to sketch the text $T$. We use random minimizers with parameters $k$ and $w=\ell-k+1$ for sketching $T$, though any type of locally-consistent sketching mechanism may be used. We start by computing the sorted list $\mathcal{M}_{\ell,k}(T)$ of minimizer positions. Let us set $z:=|\mathcal{M}_{\ell,k}(T)|$. We also consider the sequence $M[0\mathinner{.\,.}z)$ of the corresponding minimizer strings, such that $M[i]=T[p_{i}\mathinner{.\,.}p_{i}+k)$ for any position $p_{i}\in\mathcal{M}_{\ell,k}(T)$. Let $c$ be the number of distinct minimizers in $M$. We can then identify each minimizer $v\in M$ with a unique identifier (or “ID” for short, in the following) in $[0,c)$ using a map $H:\Sigma^{k}\to[0,c)$. The sketch $S[0\mathinner{.\,.}z)=\textup{{Sketch}}(T)$ is the sequence of IDs $S[i]=H(M[i])$ for all $i\in[0,z)$, which is encoded in a suitable alphabet.
@@
Two remarks about the map $H$ are in order. When $k$ is small enough to have $\sigma^{k}\ll n/\ell$, then most $k$-mers are likely to be minimizers and the map $H$ can thus be completely omitted. In what follows, we assume the case when $H$ exists. When $k$ is large, on the other hand, storing each minimizer in $M$ and the map $H$ could take a lot of space, e.g., $\mathcal{O}(c(k\log_{2}(\sigma)+\log_{2}(c)))$ bits. We can reduce the $k\log_{2}(\sigma)$ component of this space usage by first hashing the minimizers using, e.g., a rolling hash function as explained in Section 2, and only storing the mapping from minimizer hashes to their IDs. This reduces the space to $\mathcal{O}(c(q+\log_{2}(c)))$ bits, where $q$ is the number of bits used for each hash code, provided that $q<k\log_{2}(\sigma)$.
@@
Constructing the Index. Let $\Sigma^{\prime}=[0,2^{\tau})$ denote the integer alphabet that we choose to encode $S$, for some input parameter $\tau$. Further let Index denote the indexing data structure that we apply on $S$. Namely, we construct the Index of $S$, over the alphabet $\Sigma^{\prime}$, with $\tau=\log_{2}|\Sigma^{\prime}|$. Note the purpose of setting the value of $\tau$: it lets the user control the size of the alphabet we choose to encode $S$ as something that lies in $[2,n]$. Thus, we interpret each $\lceil\log_{2}(c)\rceil$-bit ID $S[i]$ as a sequence of $b:=\lceil\lceil\log_{2}(c)\rceil/\tau\rceil$ $\tau$-bit integers [^1] In the unlikely event of $z\lceil\lceil\log_{2}(c)\rceil/\tau\rceil>n$, we can either increase $\tau$ to have $z\lceil\lceil\log_{2}(c)\rceil/\tau\rceil\leq n$ or simply set $S:=T$.. This is a useful feature because some compressed full-text indexes, like the FM-index [13] or the $r$-index [14], take advantage of the repetitiveness of the text $T$ to improve its compression.
@@
Implicit Sketched Text. Note that the Index of $S$ may or may not require storing the sketched text $S$ itself. For example, the FM-index is a self-index and replaces $S$ with its compressed form. On the other hand, the suffix array is not a self-index and does require $S$. In the latter case, we can either store $S$ explicitly, or we can reconstruct $S$ on-the-fly as needed using only $T$, $H$, and $\mathcal{M}_{\ell,k}(T)$.
@@
Querying. We now describe how to compute the set $L=\textup{{Locate}}(P,T)$, given a query pattern $P[0\mathinner{.\,.}m)$ that is sufficiently long (i.e., $m\geq\ell$).
First, $P$ is sketched similarly to the text $T$, obtaining a string $Q=\textup{{Sketch}}(P)$. Specifically, its minimizer positions $\mathcal{M}_{\ell,k}(P)$ are found. Since the pattern has length $m\geq\ell$, it has at least one minimizer and we indicate with $\alpha$ and $\beta$ the position of the first and last minimizer of $P$, respectively. If one of the minimizers $P[p_{i}\mathinner{.\,.}p_{i}+k)$ of $P$, for $p_{i}\in\mathcal{M}_{\ell,k}(P)$, does not occur in the text $T$ and hence is not assigned an ID by $H$, this directly implies that $P$ does not occur in $T$. Otherwise, the list of corresponding IDs is determined as $H(P[p_{i}\mathinner{.\,.}p_{i}+k))$, for all $p_{i}\in\mathcal{M}_{\ell,k}(P)$, and this is encoded into the sketched query string $Q$ using $b$ $\tau$-bit integers per minimizer.
We locate $Q$ in $S$ using the Index of $S$. Let $L=\varnothing$ be the list of occurrences. For every position $p\in L^{\prime}=\textup{{Locate}}(S,Q)$, we first check whether $p\equiv 0\pmod{b}$. If not, the candidate match is a false positive caused by the reduction of alphabet size. Otherwise, we retrieve the position $l:=\mathcal{M}_{\ell,k}(T)[p/b]$ and verify whether $T[l-\alpha\mathinner{.\,.}l-\alpha+m)=P$ in $\mathcal{O}(m)$ time. If so, position $(l-\alpha)$ is added to $L$. Figure 2 illustrates an example.
@@
Theoretical Guarantees. We now explain how to verify an occurrence at position $p\in L^{\prime}$ in $\mathcal{O}(1)$ time and using $\mathcal{O}(z)$ space. Let the occurrence be $S[p\mathinner{.\,.}q]$ where $q=p+|Q|-1$.
@@
For faster querying in theory or for very long patterns in practice, we also store an array $F$ of fingerprints, where $F[i]=\phi(T[0\mathinner{.\,.}p_{i}])$ for all $p_{i}\in\mathcal{M}_{\ell,k}(T)$, and $\phi$ is the rolling hash function; see Section 2. This array can be constructed in $\mathcal{O}(n)$ time and has size $\mathcal{O}(z)$. Let $\mathcal{X}=\{T[p_{i}\mathinner{.\,.}p_{i}+\ell)\mid p_{i}\in\mathcal{M}_{\ell,k}(T)\}$ and $\mathcal{X}^{R}=\{T[p_{i}-\ell+1\mathinner{.\,.}p_{i}]^{R}\mid p_{i}\in\mathcal{M}_{\ell,k}(T)\}$, where $s^{R}$ denotes the reverse of the string $s$. We construct the tries $\mathcal{T}=\textup{{Trie}}(\mathcal{X})$ and $\mathcal{T}^{R}=\textup{{Trie}}(\mathcal{X}^{R})$. We label the leaf nodes representing the string $s=T[p_{i}\mathinner{.\,.}p_{i}+\ell)$ and $s^{R}=(T[p_{i}-\ell+1\mathinner{.\,.}p_{i}])^{R}$ in both tries by the set $X_{s}=\{p_{i}\mid T[p_{i}\mathinner{.\,.}p_{i}+\ell)=s\wedge p_{i}\in\mathcal{M}_{\ell,k}\}$. Each leaf is also assigned a lex-rank that is obtained via an in-order DFS traversal of the trie. We also implement an inverse function that takes $p_{i}$ as input and returns the lex-rank of the leaf node that represents $s=T[p_{i}\mathinner{.\,.}p_{i}+\ell)$ in $\mathcal{T}$. We implement the analogous inverse function for $\mathcal{T}^{R}$. Each branching node $u$ in $\mathcal{T}$ stores an interval whose left and right endpoints are the lex-rank of the leftmost and rightmost leaf node, respectively, in the subtree rooted at $u$. This information is also computed via a DFS traversal. We store the analogous information for the branching nodes in $\mathcal{T}^{R}$. Since $\mathcal{T}$ and $\mathcal{T}^{R}$ are compacted and $\sum|X_{s}|=z$, it follows that the tries and the inverse functions take $\mathcal{O}(z)$ space. The tries and the inverse functions can be constructed in $\mathcal{O}(n)$ time [7].
@@
Let us explain how these additional structures can help us verify an occurrence $S[p\mathinner{.\,.}q]$ of $Q$ in $S$ in $\mathcal{O}(1)$ time; see Figure 3. Let $l:=\mathcal{M}_{\ell,k}(T)[p/b]$ and $r:=\mathcal{M}_{\ell,k}(T)[q/b]$. Using the vector $F$, we compute $\phi(T[l+1\mathinner{.\,.}r])$ in $\mathcal{O}(1)$ time for Fact 1, because we have $F[l]=\phi(T[0\mathinner{.\,.}l])$ and $F[r]=\phi(T[0\mathinner{.\,.}r])$. We also compute the fingerprint $\phi(P[\alpha+1\mathinner{.\,.}\beta])$ once in $\mathcal{O}(m)$ time and compare the two fingerprints in $\mathcal{O}(1)$ time [^2] If $|\mathcal{M}_{\ell,k}(P)|=1$, then we always return a positive answer for this comparison.. If they are not equal, then $(l-\alpha)$ is not a valid occurrence. If they are equal, we need to check $P[0\mathinner{.\,.}\alpha]$ and $P[\beta\mathinner{.\,.}m)$. The remaining letters on each edge cannot be more than $\ell$ (by the density of the minimizers mechanism), and so the verification would cost $\mathcal{O}(\ell)$ time if we did it by letter comparisons. We can verify the edges in $\mathcal{O}(1)$ time using tries. In a preprocessing step, we spell $P[\beta\mathinner{.\,.}m)$ in $\mathcal{T}$ arriving at node $u$; and we spell $(P[0\mathinner{.\,.}\alpha])^{R}$ in $\mathcal{T}^{R}$ arriving at node $u^{\prime}$. We can then check whether $r$ is a leaf node in the subtree induced in $\mathcal{T}$ using the inverse function and the interval stored in $u$. We can also check whether $l$ is a leaf node in the subtree induced in $\mathcal{T}^{R}$ using the inverse function and the interval stored in $u^{\prime}$. This takes $\mathcal{O}(1)$ time per pair $(l,r)$. We then have that $(l-\alpha)$ is a valid occurrence if and only if both leaf nodes are located in the induced subtrees.
@@
Example. Let us now consider a practical instantiation of Theorem 4.1. Let Index be the suffix array [28] enhanced with the longest common prefix (LCP) array [25]. We choose $\tau:=\lceil\log_{2}n\rceil$ because the suffix array can be constructed in $t(n,2^{\tau})=\mathcal{O}(n)$ time, for any integer alphabet of size $2^{\tau}\leq n$, and it has size $s(n,2^{\tau})=\mathcal{O}(n)$ [23]. Given the suffix array, the LCP array can be constructed in $\mathcal{O}(n)$ time [25]. By applying Theorem 4.1, we construct a string $S$ of length $z\leq n$ over the alphabet $[0,n)$. Thus, we will construct our index in $\mathcal{O}(n+t(z,n))=\mathcal{O}(n)$ time of $\mathcal{O}(z+s(z,n))=\mathcal{O}(z)$ size. Note that, by using minimizers [31, 32], $z$ can be much smaller than $n$ in practice, for a sufficiently large value of $\ell$. Specifically, we have $z=nd$ where $d\geq 1/(\ell-k+1)$ is the density of the specific minimizer scheme used. For querying, we have $q(m,n,\sigma)=\mathcal{O}(m+\log n+\textup{{Count}}(P,T))$ when LCP information is used [28]. Thus, our query time is $\mathcal{O}(m+\log z+\textup{{Count}}(Q,S))$ because $m\geq|Q|$ and $\textup{{Count}}(Q,S)\geq\textup{{Count}}(P,T)$. Note that although $\textup{{Count}}(Q,S)\geq\textup{{Count}}(P,T)$, we also have $\log z\leq\log n$, and so beyond space savings, the resulting index can also be competitive or faster in querying.

### v3

@@ 4 A Universal Indexing Framework for Matching Long Patterns
In this section, we describe a universal indexing framework for a text $T$ of length $n$ — referred to as the U-index — to retrieve all occurrences of a pattern $P$ of length $m\geq\ell$ in $T$. Refer to Figure 1 for an overview of the proposed framework.
Figure 1: The U-index framework. Steps (1) and (2) are to build the index. The steps (3)–(6) are to query with the framework. The sketching scheme in steps (1) and (3) must be the same.
Overview.
@@
The core idea of the U-index is to sketch the text $T$. We use random minimizers with parameters $k$ and $w:=\ell-k+1$ for sketching $T$, though any type of locally consistent sketching mechanism may be used. We start by computing the sorted list $\mathcal{M}_{\ell,k}(T)$ of the minimizer positions. Let us set $z:=|\mathcal{M}_{\ell,k}(T)|$. We also consider the sequence $M[0\mathinner{.\,.}z)$ of the corresponding minimizer strings, such that $M[i]=T[p_{i}\mathinner{.\,.}p_{i}+k)$ for any position $p_{i}\in\mathcal{M}_{\ell,k}(T)$. Let $c$ be the number of distinct minimizers in $M$. We can then identify each minimizer $v\in M$ with a unique identifier (or “ID” for short, in the following) in $[0,c)$ using a map $H:\Sigma^{k}\to[0,c)$. The sketch $S[0\mathinner{.\,.}z)=\textup{{Sketch}}(T)$ is the sequence of IDs $S[i]:=H(M[i])$ for all $i\in[0,z)$, which is encoded in a suitable alphabet.
@@
Two remarks about the map $H$ are in order. When $k$ is small enough to have $\sigma^{k}\ll n/\ell$, then most $k$-mers are likely to be minimizers and the map $H$ can thus be completely omitted. In what follows, we assume the case when $H$ exists. When $k$ is large, on the other hand, storing each minimizer in $M$ and the map $H$ could take a lot of space, e.g., $\mathcal{O}(c(k\log_{2}(\sigma)+\log_{2}(c)))$ bits. We can reduce the $k\log_{2}(\sigma)$ term of this space usage by first hashing the minimizers using, e.g., a rolling hash function (see Section 2), and only storing the mapping from minimizer hashes to their IDs. This reduces the space to $\mathcal{O}(c(q+\log_{2}(c)))$ bits, where $q$ is the number of bits used for each hash code, provided that $q<k\log_{2}(\sigma)$.
Constructing the Index.
@@
Let $\Sigma^{\prime}=[0,2^{\tau})$ denote the integer alphabet that we choose to encode $S$, for some input parameter $\tau$. Further, let Index denote the chosen indexing data structure that we apply on $S$. Namely, we construct the Index of $S$, over the alphabet $\Sigma^{\prime}$, with $\tau=\log_{2}|\Sigma^{\prime}|$. Note the purpose of setting the value of $\tau$: it lets the user control the size of the alphabet we choose to encode $S$ as something that lies in $[2,n]$. Thus, we interpret each $\lceil\log_{2}(c)\rceil$-bit ID $S[i]$ as a sequence of $b:=\lceil\lceil\log_{2}(c)\rceil/\tau\rceil$ $\tau$-bit integers [^1] In the unlikely event of $z\lceil\lceil\log_{2}(c)\rceil/\tau\rceil>n$, we can either increase $\tau$ to have $z\lceil\lceil\log_{2}(c)\rceil/\tau\rceil\leq n$ or simply set $S:=T$.. This is a useful feature because some compressed full-text indexes, like the FM-index [14] or the $r$-index [15], take advantage of the repetitiveness of the text $T$ to improve its compression.
Implicit Sketched Text.
@@
Note that the Index of $S$ may or may not require storing the sketched text $S$ itself. For example, the FM-index is a self-index and replaces $S$ with its compressed form. On the other hand, the suffix array is not a self-index and does require $S$. In the latter case, we can either store $S$ explicitly, or we can reconstruct $S$ on-the-fly as needed using only $T$, $H$, and $\mathcal{M}_{\ell,k}(T)$.
@@
Querying.
We now describe how to compute the set $L=\textup{{Locate}}(P,T)$, given a query pattern $P[0\mathinner{.\,.}m)$ that is sufficiently long (i.e., $m\geq\ell$).
First, $P$ is sketched similarly to the text $T$, obtaining a string $Q=\textup{{Sketch}}(P)$. Specifically, its minimizer positions $\mathcal{M}_{\ell,k}(P)$ are found. Since the pattern has length $m\geq\ell$, it has at least one minimizer, and we indicate with $\alpha$ and $\beta$ the position of the first and last minimizer of $P$, respectively. If one of the minimizers $P[p_{i}\mathinner{.\,.}p_{i}+k)$ of $P$, for $p_{i}\in\mathcal{M}_{\ell,k}(P)$, does not occur in the text $T$ and hence is not assigned an ID by $H$, this directly implies that $P$ does not occur in $T$. Otherwise, the list of corresponding IDs is determined as $H(P[p_{i}\mathinner{.\,.}p_{i}+k))$, for all $p_{i}\in\mathcal{M}_{\ell,k}(P)$, and this is encoded into the sketched query string $Q$ using $b$ $\tau$-bit integers per minimizer.
We locate $Q$ in $S$ using the Index of $S$. Let $L=\varnothing$ be the initially empty list of occurrences. For every position $p\in L^{\prime}=\textup{{Locate}}(S,Q)$, we first check whether $p\equiv 0\pmod{b}$. If not, the candidate match is a false positive caused by the reduction of alphabet size. Otherwise, we retrieve the position $l:=\mathcal{M}_{\ell,k}(T)[p/b]$ and verify whether $T[l-\alpha\mathinner{.\,.}l-\alpha+m)=P$ in $\mathcal{O}(m)$ time. If so, position $(l-\alpha)$ is added to $L$. Figure 2 illustrates an example.
@@
We now explain how to verify an occurrence at position $p\in L^{\prime}$ in $\mathcal{O}(1)$ time using $\mathcal{O}(z)$ space. Let the occurrence be $S[p\mathinner{.\,.}q]$, where $q=p+|Q|-1$.
@@
For faster querying in theory or for very long patterns in practice, we also store an array $F$ of fingerprints, where $F[i]=\phi(T[0\mathinner{.\,.}p_{i}])$, for all $p_{i}\in\mathcal{M}_{\ell,k}(T)$, and $\phi$ is the rolling hash function; see Section 2. This array can be constructed in $\mathcal{O}(n)$ time and has size $\mathcal{O}(z)$. Recall that $\ell$ is the lower bound on the length of $P$. Let $\mathcal{X}=\{T[p_{i}\mathinner{.\,.}p_{i}+\ell)\mid p_{i}\in\mathcal{M}_{\ell,k}(T)\}$ and $\mathcal{X}^{R}=\{(T[p_{i}-\ell+1\mathinner{.\,.}p_{i}])^{R}\mid p_{i}\in\mathcal{M}_{\ell,k}(T)\}$, where $s^{R}$ denotes the reverse of the string $s$. We construct the tries $\mathcal{T}:=\textup{{Trie}}(\mathcal{X})$ and $\mathcal{T}^{R}:=\textup{{Trie}}(\mathcal{X}^{R})$. We label the leaf nodes representing the string $s=T[p_{i}\mathinner{.\,.}p_{i}+\ell)$ and $s^{R}=(T[p_{i}-\ell+1\mathinner{.\,.}p_{i}])^{R}$ in both tries by the set $X_{s}=\{p_{i}\mid T[p_{i}\mathinner{.\,.}p_{i}+\ell)=s\wedge p_{i}\in\mathcal{M}_{\ell,k}\}$. Each leaf node is also assigned a lex-rank that is obtained via an in-order DFS traversal of the trie. We also implement an inverse function that takes $p_{i}$ as input and returns the lex-rank of the leaf node that represents $s=T[p_{i}\mathinner{.\,.}p_{i}+\ell)$ in $\mathcal{T}$. We implement the analogous inverse function for $\mathcal{T}^{R}$. Each branching node $u$ in $\mathcal{T}$ stores an interval whose left and right endpoints are the lex-rank of the leftmost and rightmost leaf node, respectively, in the subtree rooted at $u$. This information is also computed via a DFS traversal. We store the analogous information for the branching nodes in $\mathcal{T}^{R}$. Since $\mathcal{T}$ and $\mathcal{T}^{R}$ are compacted and $\sum|X_{s}|=z$, it follows that the tries and the inverse functions take $\mathcal{O}(z)$ space. Furthermore, they can be constructed in $\mathcal{O}(n)$ time [8].
@@
Let us explain how these additional structures can help us verify an occurrence $S[p\mathinner{.\,.}q]$ of $Q$ in $S$ in $\mathcal{O}(1)$ time; see Figure 3. Let $l:=\mathcal{M}_{\ell,k}(T)[p/b]$ and $r:=\mathcal{M}_{\ell,k}(T)[q/b]$. Using the vector $F$, we compute $\phi(T[l+1\mathinner{.\,.}r])$ in $\mathcal{O}(1)$ time by Fact 1, because we have $F[p]=\phi(T[0\mathinner{.\,.}l])$ and $F[q]=\phi(T[0\mathinner{.\,.}r])$. We also compute the fingerprint $\phi(P[\alpha+1\mathinner{.\,.}\beta])$ once in $\mathcal{O}(m)$ time and compare the two fingerprints in $\mathcal{O}(1)$ time [^2] If $|\mathcal{M}_{\ell,k}(P)|=1$, then we always return a positive answer for this comparison.. If they are not equal, then $(l-\alpha)$ is not a valid occurrence. If they are equal, we need to check $P[0\mathinner{.\,.}\alpha]$ and $P[\beta\mathinner{.\,.}m)$. The remaining letters on each edge cannot be more than $\ell$ (by the density of the minimizers mechanism), and so the verification would cost $\mathcal{O}(\ell)$ time if we did it by letter comparisons. We can verify the edges in $\mathcal{O}(1)$ time using tries; see Figure 3. In a preprocessing step, we spell $P[\beta\mathinner{.\,.}m)$ in $\mathcal{T}$ arriving at node $u$; and we spell $(P[0\mathinner{.\,.}\alpha])^{R}$ in $\mathcal{T}^{R}$ arriving at node $u^{\prime}$. This takes $\mathcal{O}(\ell)$ time. We can then check whether $r$ is a leaf node in the subtree induced in $\mathcal{T}$ using the inverse function and the interval stored in $u$. We can also check whether $l$ is a leaf node in the subtree induced in $\mathcal{T}^{R}$ using the inverse function and the interval stored in $u^{\prime}$. This takes $\mathcal{O}(1)$ time per pair $(l,r)$. We then have that $(l-\alpha)$ is a valid occurrence if and only if both leaf nodes are located in the induced subtrees.
@@
Example.
Let us now consider a practical instantiation of Theorem 4.1. Let Index be the suffix array [31] enhanced with the longest common prefix (LCP) array [26]. We choose $\tau:=\lceil\log_{2}n\rceil$ because the suffix array can be constructed in $t(n,2^{\tau})=\mathcal{O}(n)$ time, for any integer alphabet of size $2^{\tau}\leq n$, and it has size $s(n,2^{\tau})=\mathcal{O}(n)$ [24]. Given the suffix array, the LCP array can be constructed in $\mathcal{O}(n)$ time [26]. By applying Theorem 4.1, we construct a string $S$ of length $z\leq n$ over the alphabet $[0,n)$. Thus, we will construct our index of $\mathcal{O}(z+s(z,n))=\mathcal{O}(z)$ size in $\mathcal{O}(n+t(z,n))=\mathcal{O}(n)$ time. Note that, by using minimizers [34, 35], $z$ can be much smaller than $n$ in practice, for a sufficiently large value of $\ell$. Specifically, we have $z=nd$ where $d\geq 1/(\ell-k+1)$ is the density of the specific minimizer scheme used. For querying, we have $q(m,n,2^{\tau})=\mathcal{O}(m+\log n+\textup{{Count}}(P,T))$ when LCP information is used [31]. Thus, our query time is $\mathcal{O}(m+\log z+\textup{{Count}}(Q,S))$ because $m\geq|Q|$ and $\textup{{Count}}(Q,S)\geq\textup{{Count}}(P,T)$. We stress that even if $\textup{{Count}}(Q,S)\geq\textup{{Count}}(P,T)$, we also have $\log z\leq\log n$, and so beyond space savings, the resulting index can also be competitive or faster in querying.